After resolving the vectors into their components
Answer #1

Consider the forces on the pith balls.

The Forces are m*g down and Electrostatic Force horizontally.

Since the string is at theta° and the forces act at right angles

F_electro / F_gravity or E_electro = kq2/r2 = m*g*tantheta°

q = sqrt ( m*g*tantheta° *r^2/k)